Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

combinations. Auto repairmen

Expert replies
by gmatruler » Tue Jul 13, 2010 5:05 am
Joshua and jose work at an auto repair center with 4 other workers. For a survey on healthcare insurance, 2 of the 6 workers will be randomly chosen to be interviewed. What is the probability that Joshua and Jose will both be chosen?

A) 1/15
B) 1/12
C) 1/9
D) 1/6
E) 1/3
Join the discussion
Source: — Problem Solving |

by Patrick_GMATFix » Tue Jul 13, 2010 5:10 am
This is a combination problem. There are 6 people from which 2 will be selected. This means that the total number of ways to pick people is a combination of 2 from 6: 6!/(2!4!) = 15.

Of these 15 possible group, only one group of 2 will include both Joshua and Jose. Therefore the probability that they are both picked is 1/15

The correct answer is A. A more detailed explanation as well as video solution can be viewed at GMATPrep Question 1208. To practice similar questions, set topic='combinatorics' and difficulty='600-700 & 700+' in the Drill Generator

Good luck,
-Patrick
  • Ask me about tutoring.
Join the discussion

by Rahul@gurome » Tue Jul 13, 2010 5:19 am
2 out of 6 workers are to be chosen randomly, there are 6C2 ways of doing this, out of which there is just one possibility that Joshua and Jose will both be chosen.
6C2 = 6!/2!(6-2)! = 6!/(2!4!)= 15

Out of these 15 pairs just one pair will be that of Joshua and Jose.
So, required probability = 1/15

The correct answer is (A).
Rahul Lakhani
Quant Expert
Gurome, Inc.
https://www.GuroMe.com
On MBA sabbatical (at ISB) for 2011-12 - will stay active as time permits
1-800-566-4043 (USA)
+91-99201 32411 (India)
Join the discussion

by GMATGuruNY » Tue Jul 13, 2010 7:55 am
gmatruler wrote:Joshua and jose work at an auto repair center with 4 other workers. For a survey on healthcare insurance, 2 of the 6 workers will be randomly chosen to be interviewed. What is the probability that Joshua and Jose will both be chosen?

A) 1/15
B) 1/12
C) 1/9
D) 1/6
E) 1/3
We also could handle this as a straight probability question.

Probability = (good outcomes)/(total outcomes)

In this problem:

good outcome = Josh or Jose
total outcomes = total number of people


Probability of picking Josh or Jose on the first pick: 2/6 (because there are 2 good outcomes out of a total of 6 people)

Probability of picking Josh or Jose on the second pick: 1/5 (because if Josh or Jose was chosen on the first pick, then only 1 good outcome is left for the 2nd pick, out of a total of 5 remaining people)

We need both of these events to happen. When we want multiple events to happen, we multiply the fractions:

P(A and B) = P(A) * P(B)

We multiply the fractions because the more events we want to happen together, the smaller the probability, and when we multiply fractions, the result just keeps getting smaller.

2/6 * 1/5 = 2/30 = 1/15.

The correct answer is A.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by gmatruler » Sun Jul 18, 2010 6:37 am
Wow 3 different, expert ways to solve. Thank you :-)
Join the discussion